The least distance between the circles |z| =12 and |z-3-4i| = 5 is....Explain the steps.
 

Grade:

1 Answers

Arun
5 years ago
Dear student
 
|z| = 12 means a circle of radius 12 with centre at origin (0 , 0)
 
Now
|z – 3 – 4i| = 5 means a circle of radius 5 wih centre at (3, 4)
 
Now for least distance
 
1 2 – 5 – |3 + 4i| = 7 – 5 = 2 units